OpenStudy (anonymous):

Raul calculated that he would spend $125 on school supplies this year. He actually spent $87.50 on school supplies. What is Raul’s percent of error? 42.9% 37.5% 30% 3%

OpenStudy (anonymous):

@itsbribro

OpenStudy (itsbribro):

a. 42.9% :) Medal?

OpenStudy (anonymous):

ok :)

OpenStudy (itsbribro):

yep :)

OpenStudy (ali2x2):

So you would use: [ ((V2 - V1) / |V1|) * 100 ] V1 = 87.5 to V2 = 125 = ((125 - 87.5) / |87.5|) * 100 = (37.5 / 87.5) * 100 = 0.428571 * 100 = 42.8571% change = 42.8571% increase

OpenStudy (ali2x2):

So in this case, shes right.